# Massive Discovery of Low-Dimensional Materials from Universal Computational Strategy

### v1

@@ III Discussion
We demonstrated that both MACE and MatterSim predict interatomic force constants and material dimensionalities with an accuracy comparable to first-principles calculations. Using the slightly better-performing MatterSim, we calculated phonons for 35,68935{,}689 Materials Project structures, kept the stable ones, and classified their dimensionalities using a method not based on geometries but on true chemical bonding between the atoms. Consequently, we discovered $9139$ low-dimensional materials: $1838$ clusters, $1760$ chains, $3057$ sheets, and $2484$ mixed-dimensional systems. For the discovered 2D materials, we further calculated the interlayer binding energies and found that $152$ are easily exfoliable and $808$ are potentially exfoliable 2D materials. These 2D materials do not exist in the known 2D materials databases. We believe that combining ML with the bond strength-based dimensionality classification method will continue to help discover and explore low-dimensional materials for future applications. We also hope that our discovery inspires experimental work and leads to further expansion of the low-dimensional materials catalog.
IV Methods
Computational methods
The machine-learning (ML) calculations were performed using MatterSim with pre-trained model v1.0.0-5M Yang et al. (2024) and MACE Batatia et al. (2022a, b) with the pre-trained large model of MACE-MP-0 (MACE-MP-0b3) without dispersion. The MatterSim-v1 model is based on the M3GNet Chen and Ong (2022) architecture, which was trained on relaxations obtained from the Materials Project Jain et al. (2013). Both UMLIP models are available as calculators in the Atomic Simulation Environment (ASE) Larsen et al. (2017). ASE was used, with the help of Python, to construct the calculation workflow. In the benchmark, all parameters were kept the same as those in the Phonon database Togo and Tanaka (2015). All new materials were relaxed using the FIRE Bitzek et al. (2006) and BFGS algorithm to forces below $1$ meV/Å. The crystal structures and their properties were extracted using MP-API Ong et al. (2015) and Python materials genomics (Pymatgen) Ong et al. (2013). The space groups are calculated using spglib Atsushi Togo and Tanaka (2024) with symmetry precision $10^{-6}$.
Initial dimensionalities were calculated using Larsen Larsen et al. (2019) method, which is implemented in ASE Larsen et al. (2017). Phonon calculations were done using the Phonopy package Togo et al. (2023); Togo (2023), using a displacement of $0.01$ Å. The supercell sizes used for determining the force constants were chosen so that the number of atoms was less than or equal to $250$ while keeping the point groups of the Bravais lattices of the conventional unit cell and the supercell the same.
Initial structures of known 2D materials (Fig. 4a) were extracted from Ref. 27. All density-functional theory calculations were performed using VASP (Vienna Ab initio Simulation Package) software Kresse and Furthmüller (1996, 1996), together with the projector-augmented wave method Blöchl (1994). The Becke-Johnson (BJ) damping variant of the DFT-D3 functional was adopted Grimme et al. (2011). The plane wave cutoff is set to $520$ eV. The Brillouin zone of the unit cell is sampled by a $\Gamma$-centered k-point mesh whose density is defined by $R_{k}=20$, which in VASP determines the subdivisions $N_{1}$, $N_{2}$, and $N_{3}$ along the reciprocal lattice vectors $b_{1}$, $b_{2}$, and $b_{3}$, respectively, via $N_{i}={\rm max}(1,R_{k}|b_{i}|+0.5)$ and rounded to an integer. The energy convergence criterion between consequential steps is set to $10^{-8}$ eV for the self-consistent loop and $10^{-6}$ eV for ionic relaxation. Structures were considered to be relaxed when no force exceeded $1$ meV/Å.
In the machine-learning binding energy calculation, we employed torch-dfd3 Grimme (2006); Grimme et al. (2010, 2011); Takamoto et al. (2022) alongside MatterSim to add the D3 correction in our calculations to correct the weak vdW interaction.
Binding energy calculation
The method used to obtain the binding energy $E_{b}$ is adapted from Ref. 24. Taking materials from MP, the bulk unit cell is first relaxed with D3 correction to obtain the initial energy $E_{0}$ as well as the force constant. Using the latter, 2D units are identified using FCDimen. The resulting monolayers are then isolated and relaxed, leading to the layer energies $E_{l}$. The binding energy $E_{b}$ is then found as
$E_{b}=\frac{\sum E_{l}-E_{0}}{A\cdot N_{l}},$ (6)
@@ IV Methods
Acknowledgments
We acknowledge the Jane and Aatos Erkko Foundation for funding (project EcoMet) and the Research Council of Finland for support under Academy Project No. 357483. We also thank CSC–IT Center for Science Ltd. for generous grants of computer time.

### v2

@@ Conclusion
We demonstrated that both MACE and MatterSim predict interatomic force constants and material dimensionalities with an accuracy comparable to first-principles calculations. Using the slightly better-performing MatterSim, we calculated phonons for 35,68935{,}689 Materials Project structures, kept the stable ones, and classified their dimensionalities using a method not based on geometries but on true chemical bonding between the atoms. Consequently, we discovered $9139$ low-dimensional materials: $1838$ clusters, $1760$ chains, $3057$ sheets, and $2484$ mixed-dimensional systems. For the discovered 2D materials, we further calculated the interlayer binding energies and found that $146$ are easily exfoliable and $741$ are potentially exfoliable 2D materials. These 2D materials do not exist in the known 2D materials databases. We believe that combining ML with the bond strength-based dimensionality classification method will continue to help discover and explore low-dimensional materials for future applications. We also hope that our discovery inspires experimental work and leads to further expansion of the low-dimensional materials catalog.
@@
Supporting information
The following file is available free of charge.
  • Supporting Information: Additional details on benchmark results of universal machine-learning interatomic potentials, a list of found low-dimensional materials with the highest score, sample Python scripts to facilitate the usage of databases, and phonon dispersion of selected 2D materials. (PDF)
Acknowledgements
We acknowledge the Jane and Aatos Erkko Foundation for funding (project EcoMet) and the Research Council of Finland for support under Academy Project No. 357483. The authors thank the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call. We also thank CSC–IT Center for Science Ltd. for generous additional grants of computer time.